SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (August 15, 2006)

Attendees: Debbie Bard, Rainer Bartoldus, Al Eisner, Selina Li, Rahmat Rahmat, Dong Su, Mark Tibbetts

Operations

  • BaBar/PEP shutdown will begin next Monday around 6am. Two items that we know for sure we would like to test before the end of run5b.
    • a new trigger configuration with ZPD configuration change (putting in the 1Zn to replace the 1Zk for the 1Z trigger line): Rainer is working on getting the configration ready to test at IR2.
    • L3 tracking code test (to pick up more L3 tracks, etc.): Rainer and Arik are working on this.

EMT

  • Dan is on vacation, but will be back next week to show Mark more EMT stuff. Mark will take over Dan's full responsibility for EMT in September when Dan heads back to UK.

DCT

  • Su Dong and Selina used the oscilloscope to look at the frame bit, clk and TSF data at the backplane of the TSf board at the teststand. Nothing really stands out when compared the old firmare to the version which caused ZPD damage even though the TSF data seem fine. Not sure if Marc will be able to produce any version of the TSF firmware to test before the shutdown next week.

Data Quality and Monitoring

  • Rahmat showed the plots for lumi ratio. He confirms that the labels on the plots are correct, but we prefer to have B counting mu-pairs as the denominator when calculating the lumi ratio. Rahmat will get this revised and also work on using time on the x-axis.
  • After the new online release installed at IR2 last Tuesday, the fastmon .ps archive did not have from runs 67434-67614 (inclusive). In particular, the EMT efficiency plots were not produced correctly. Jim Hamilton found that there was something wrong with the online release that made the fastmon build crash. He built again from 18.7.0 without online and added in both L1TFMon and L1TOepSequences, and it worked.
  • Su Dong showed some T0 plots for a good run 60949 and a high lumi and high background run 67092. He was surprised that background got into the T0 plots, particularly the bhabha and mu-pair. He suggested Mark can probably help investigate these events with T0 < 200ns in the bhabha plot to see what they really are and who triggered them.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.